OpenStudy (anonymous):

if sin(theta) = 3/5 how do i find the exact value of sin3(theta)

OpenStudy (anonymous):

Are you trying to solve for θ?

OpenStudy (dumbcow):

http://en.wikipedia.org/wiki/Trig_Identity#Double-angle.2C_triple-angle.2C_and_half-angle_formulae \[\sin 3 \theta = -4 (\frac{3}{5})^3 + 3(\frac{3}{5})\]

OpenStudy (dls):

\[\Huge \sin 3 \theta = 3 \sin \theta-4 \sin^3 \theta\]

OpenStudy (anonymous):

\[\sin ^3\left(\sin ^{-1}\left(\frac{3}{5}\right)\right)=\frac{27}{125} \]